OpenStudy (anonymous):

(2,9),x+7y=4 Write an equation of the line containing the given point and parallel to the given line. Express your answer in the given form y=mx+b

OpenStudy (karatechopper):

y-9=1/7(x-2) y-9=1/7x-2/7 y-9+9=1/7x-2/7+9 y=1/7x+ 8 5/7

OpenStudy (karatechopper):

or u can write it like y=1/7x+61/7